# Compatible $E$-differential forms on Lie algebroids over (pre-)multisymplectic manifolds

### v3

@@ 4 Examples
$(\mathrm{d}^{\nabla}+{}^{E}\mathrm{d}^{\nabla})\mu=-\sum_{k=0}^{n}\iota_{\rho}^{n+1-k}{\omega}.$ (42)
Expanding the equation by form degree of both sides, Equation (42) is the following $n$ equations,
$\mathrm{d}^{\nabla}\mu_{n-1}=-\iota_{\rho}^{1}{\omega},$ (43)
$\mathrm{d}^{\nabla}\mu_{n-2}+{}^{E}\mathrm{d}^{\nabla}\mu_{n-1}=-\iota_{\rho}^{2}{\omega},$ (44)
@@
$\mathrm{d}^{\nabla}\mu_{k-1}+{}^{E}\mathrm{d}^{\nabla}\mu_{k}=-\iota_{\rho}^{n+1-k}{\omega},$ (45)
@@
${}^{E}\mathrm{d}^{\nabla}\mu_{0}={}^{E}\mathrm{d}\mu_{0}=-\iota_{\rho}^{n+1}{\omega}.$ (46)
The final equation (46) is the condition that $J=\mu_{0}$ is a compatible $E$-$n$-form (11).
Proposition 4.9
@@
5 Higher Dirac structures on $E\oplus\wedge^{n-1}E^{*}$
@@
Let $(E,\rho,[-,-])$ be a Lie algebroid over $M$. Then, an almost Lie algebroid structure is induced on $E\oplus\wedge^{n-1}E^{*}$. Three operations are constructed from the Lie algebroid operations. They are a bilinear symmetric form ${({{-},{-}})}:\Gamma(E\oplus\wedge^{n-1}E^{*})\times\Gamma(E\oplus\wedge^{n-1}E^{*})\rightarrow\Gamma(\wedge^{n-2}E^{*})$, a bundle map ${\widehat{\rho}}:E\oplus\wedge^{n-1}E^{*}\rightarrow TM$ called the anchor map, and a bilinear bracket ${[{{-},{-}}]_{D}}:\Gamma(E\oplus\wedge^{n-1}E^{*})\times\Gamma(E\oplus\wedge^{n-1}E^{*})\rightarrow\Gamma(E\oplus\wedge^{n-1}E^{*})$ called a (higher) Dorfman bracket. Let $u+\alpha,v+\beta\in\Gamma(E\oplus\wedge^{n-1}E^{*})$, where $u,v\in\Gamma(E)$ and $\alpha,\beta\in\Gamma(\wedge^{n-1}E^{*})$. Then three operations are defined by
${({{u+\alpha},{v+\beta}})}$ $=\iota_{u}\beta+\iota_{v}\alpha,$ (47)
${\widehat{\rho}}(e)f$ $=\rho(u)f,$ (48)
${[{{u+\alpha},{v+\beta}}]_{D}}$ $=[u,v]+{}^{E}{\mathcal{L}}_{u}\beta-\iota_{v}{}^{E}\mathrm{d}\alpha+\iota_{u}\iota_{v}(\iota_{\rho}^{n+1}\omega),$ (49)
where the interior product $\iota_{v}$ is the contraction of an element $v\in\Gamma(E)$ with $\Gamma(E^{*})$, $\rho(u)$ is the anchor map of $E$, and ${}^{E}{\mathcal{L}}_{u}:=\iota_{u}{}^{E}\mathrm{d}+{}^{E}\mathrm{d}\iota_{u}$ is the $E$-Lie derivative. The Dorfman bracket (49) is not skew-symmetric but satisfies the Jacobi type identity,
${[{{{[{{u+\alpha},{v+\beta}}]_{D}}},{w+\gamma}}]_{D}}+{[{{{[{{v+\beta},{w+\gamma}}]_{D}}},{u+\alpha}}]_{D}}+{[{{{[{{w+\gamma},{u+\alpha}}]_{D}}},{v+\beta}}]_{D}}=0.$ (50)
An algebroid satisfying Equation (50) is called a Leibniz algebroid.
Especially, we call $(E\oplus\wedge^{n-1}E^{*},{({{-},{-}})},\rho,{[{{-},{-}}]_{D}})$ a Lie $n$-algebroid, or a Vinogradov algebroid induced from a Lie algebroid $E$ [20]. We only consider this specific Lie $n$-algebroid in this paper. There are many analysis about general theories for Lie $n$-algebroids. In general, Lie $n$-algebroids are induced from a QP-manifold of degree $n$. A general concept of Lie $n$-algebroids is referred to, for instance, [7, 28, 15, 16, 21, 39, 41, 45].
Equations (47)–(49) does not contain $J$. The compatible condition (11), ${}^{E}\mathrm{d}J=-\iota_{\rho}^{n+1}\omega$, is described as an extra structure on the Lie $n$-algebroid, a higher Dirac structure[8, 4, 45].
Definition 5.1
@@
$\Gamma(L)$ =$=\{u+{({{J},{u}})}\in\Gamma(E\oplus\wedge^{n-1}E^{*})|u\in\Gamma(E)\}.$ (51)
Theorem 5.2
@@
Moreover, using Eq. (49), the Dorfman bracket is computed as
@@
$=[u,v]+{}^{E}{\mathcal{L}}_{u}\iota_{v}J-\iota_{v}{}^{E}\mathrm{d}\iota_{u}J+\iota_{u}\iota_{v}(\iota_{\rho}^{n+1}\omega)$
$=[u,v]+{}^{E}{\mathcal{L}}_{u}\iota_{v}J-\iota_{v}{}^{E}{\mathcal{L}}_{u}J-\iota_{v}\iota_{u}{}^{E}\mathrm{d}J+\iota_{u}\iota_{v}(\iota_{\rho}^{n+1}\omega),$ (53)
Using the formula ${}^{E}{\mathcal{L}}_{u}\iota_{v}-\iota_{v}{}^{E}{\mathcal{L}}_{u}=\iota_{[u,v]}$ analogous to the Cartan formula, Eq. (53) becomes
$=[u,v]+{({{J},{[u,v]}})}+\iota_{u}\iota_{v}{}^{E}\mathrm{d}J+\iota_{u}\iota_{v}(\iota_{\rho}^{n+1}\omega),$ (54)
which is an element of $\Gamma(L)$ again if the identity ${}^{E}\mathrm{d}J=-\iota_{\rho}^{n+1}\omega$ is satisfied. Note that we use $\iota_{u}\iota_{v}{}^{E}\mathrm{d}J=-\iota_{v}\iota_{u}{}^{E}\mathrm{d}J$. Since then, Eq. (54) becomes
${[{{u+{({{J},{u}})}},{v+{({{J},{v}})}}}]_{D}}$ $=[u,v]+{({{J},{[u,v]}})},$ (55)
@@
${({{\alpha+u},{\beta+v}})}$ $=\iota_{u}\beta+\iota_{v}\alpha,$ (56)
${\widehat{\rho}}(\alpha+u)f$ $=\pi^{\sharp}(\alpha)(f),$ (57)
${[{{\alpha+u},{\beta+v}}]_{D}}$ $=[\alpha,\beta]_{\pi,H}+{\mathcal{L}}^{\pi,H}_{\alpha}v-\iota_{\beta}\mathrm{d}_{\pi,H}u+\langle\otimes^{3}\pi,\,H\rangle(\alpha,\beta).$ (58)
@@
$[\alpha,\beta]_{\pi,H}$ $=$ ${\mathcal{L}}_{\pi^{\sharp}(\alpha)}\beta-{\mathcal{L}}_{\pi^{\sharp}(\beta)}\alpha-\mathrm{d}(\pi(\alpha,\beta))+\iota_{\pi^{\sharp}(\alpha)}\iota_{\pi^{\sharp}(\beta)}H,$ (59)
$\mathrm{d}_{\pi}u$ $=$ $[\pi,u]_{S},$ (60)
$\mathrm{d}_{\pi,H}u$ $=$ $\mathrm{d}_{\pi}u+\langle(\otimes^{2}\pi\otimes 1)H,\,u\rangle,$ (61)
$L^{\pi,H}_{\alpha}v$ $=$ $(\iota_{\alpha}\mathrm{d}_{\pi,H}+\mathrm{d}_{\pi,H}\iota_{\alpha})v.$ (62)
Let us introduce a bivector field $J\in\Gamma(\wedge^{2}TM)$ and define a subbundle defined by
$L$ =$=\{\alpha+{({{J},{\alpha}})}\in\Gamma(T^{*}M\oplus TM)|\alpha\in\Omega^{1}(M)\}.$ (63)
Iff $L$ is a Dirac structure, $J$ is a compatible $T^{*}M$-$2$-form, i.e., a bivector field satisfying ${}^{E}\mathrm{d}J=-\iota_{\pi}^{2}H$. This case is special. We can take $J=\pi$ since both are elements of $\Gamma(\wedge^{2}TM)$, and ${}^{E}\mathrm{d}\pi=-\iota_{\pi}^{2}H$ is equivalent to the condition that $(\pi,H)$ is a twisted Poisson structure.
@@ 6 Q-manifold description
Definition 6.1
If an N-manifold ${\mathcal{M}}$ has a vector field $Q$ of degree $+1$ satisfying $Q^{2}=0$, it is called a Q-manifold, or a differential graded (dg) manifold.
@@
$Q$ $=\rho^{i}_{a}(x)q^{a}\frac{\partial}{\partial x^{i}}-\frac{1}{2}C_{bc}^{a}(x)q^{b}q^{c}\frac{\partial}{\partial q^{a}},$ (64)
where $\rho^{i}_{a}(x)$ and $C^{a}_{bc}(x)$ are local functions. We define a bundle map $\rho:E\rightarrow TM$ and the Lie bracket $[-,-]:\Gamma(E)\times\Gamma(E)\rightarrow\Gamma(E)$ as
$\rho(e_{a})$ $:=\rho^{i}_{a}(x)\partial_{i},$ (65)
$~{}[e_{a},e_{b}]$ $:=C_{ab}^{c}(x)e_{c}.$ (66)
for the basis of sections $e_{a}\in\Gamma(E)$. Then, $Q^{2}=0$ is equivalent that the two operations $(\rho,[-,-])$ satisfy the definition of a Lie algebroid on $E$: all the identities of the anchor map and the Lie bracket in the Lie algebroid $E$ are obtained from the $Q^{2}=0$ condition.
@@
Definition 6.2
If an N-manifold ${\mathcal{M}}$ has a graded symplectic form ${\omega}_{grad}$ of degree $n$ and a vector field $Q$ of degree $+1$ satisfies $Q^{2}=0$, and ${\mathcal{L}}_{Q}{\omega}_{grad}=0$, $({\mathcal{M}},{\omega}_{grad},Q)$ is called a QP-manifold of degree $n$.
In a QP-manifold, if $n\neq 0$, there exist a Hamiltonian function $\Theta$ such that $Q={\{{{\Theta},{-}}\}}$. $Q^{2}=0$ is equivalent to ${\{{{\Theta},{\Theta}}\}}=0$.
For a Lie algebroid, we take an $n-1$-shifted cotangent bundle ${\mathcal{M}}=T^{*}[n-1]E[1]$ with a nonnegative integer $n-1$. Take local coordinates on $T^{*}[n-1]E[1]$, $(x^{i},q^{a},z_{i},y_{a})$ of degree $(0,1,n-1,n-2)$. Since $T^{*}[n-1]E[1]$ is a (graded) cotangent bundle, it has a canonical graded symplectic form of degree $n-1$,
${\omega}_{grad}$ $=\delta x^{i}\wedge\delta z_{i}+\delta q^{a}\wedge\delta y_{a},$ (67)
where $\delta$ is the graded de Rham differential on the graded manifold. A graded Poisson bracket for ${\omega}_{grad}$ is defined by
${\{{{f},{g}}\}}$ $=$ $(-1)^{|f|+n}\iota_{X_{f}}\iota_{X_{g}}{\omega}_{grad}.$ (68)
@@
$\{x^{i},z_{j}\}$ $=\delta^{i}_{j},\qquad\{z_{j},x^{i}\}=-\delta^{i}_{j},$ (69)
$\{q^{a},y_{b}\}$ $=\delta^{a}_{b},\qquad\{y_{b},q^{a}\}=-(-1)^{n-2}\delta^{a}_{b}.$ (70)
@@
$\Theta_{Lie}$ $=\langle\rho(q),\,z\rangle+\frac{(-1)^{n-1}}{2}{({{[q,q]},{y}})}=(-1)^{n-1}\rho^{i}_{a}(x)z_{i}q^{a}+\frac{(-1)^{n-1}}{2}C_{bc}^{a}(x)q^{b}q^{c}y_{a}.$ (71)
Here $\langle$-,\,$\rangle$ - , - ⟩ is the pairing of $TM$ and $T^{*}M$ an (,){({{-},{-}})}( - , - ) is the pairing of $E$ and $E^{*}$. $\{\Theta_{Lie},\Theta_{Lie}\}=0$ is equivalent to $\rho(e_{a})=\rho^{i}_{a}(x)\partial_{i}$ and $C_{bc}^{a}(x)$ are the anchor map and the structure function of the Lie bracket of a Lie algebroid.
We can prove that $\iota_{Q_{Lie}}{\omega}_{grad}=-\delta\Theta_{Lie}$, equivalently, $Q_{Lie}=-\{\Theta_{Lie},-\}$ in (71) satisfies $Q_{Lie}^{2}=0$ and ${\mathcal{L}}_{Q_{Lie}}{\omega}_{grad}=0$ if and only if $E$ is a Lie algebroid. The local coordinate expression is
${Q_{\mathrm{Lie}}}$ =$=-\{\Theta_{Lie},-\}$
$=\rho^{i}_{a}(x)q^{a}\frac{\partial}{\partial x^{i}}-\frac{1}{2}C_{bc}^{a}(x)q^{b}q^{c}\frac{\partial}{\partial q^{a}}+\left((-1)^{n}\rho^{i}_{a}z_{i}-C_{ab}^{c}(x)q^{b}y_{c}\right)\frac{\partial}{\partial y_{a}}$
$-\left(\partial_{i}\rho^{j}_{a}z_{j}q^{a}-\frac{(-1)^{n}}{2}\partial_{i}C_{bc}^{a}(x)q^{b}q^{c}y_{a}\right)\frac{\partial}{\partial z_{i}}.$ (72)
Here $\partial_{i}=\frac{\partial}{\partial x^{i}}$. ${Q_{\mathrm{Lie}}}$ is also denoted by ${\mathcal{L}}_{Q}={Q_{\mathrm{Lie}}}$ and called the tangent lift of the vector field $Q$ in (64). ${Q_{\mathrm{Lie}}}$ is the homological vector field such that $Q={Q_{\mathrm{Lie}}}|_{E[1]}$. $T^{*}[n-1]E[1]$ has a QP-manifold structure induced from a Lie algebroid structure on $E$.
@@
${\underline{J}}{}$ $:=\frac{1}{n!}J_{a_{1}\ldots a_{n}}(x)q^{a_{1}}\ldots q^{a_{n}},$ (73)
@@
$\frac{1}{n!}J_{a_{1}\ldots a_{n}}(x)$ $:=J(e_{a_{1}},\cdots,e_{a_{n}}).$ (74)
The term ${\underline{J}}{}$ is summed to $\Theta_{Lie}$ as
$\Theta_{LieJ}$ $=\Theta_{Lie}+(-1)^{n-1}{\underline{J}}{},$ (75)
and consider the vector field of degree $+1$
$Q_{LieJ}$ =$=-{\{{{\Theta_{LieJ}},{-}}\}}.$ (76)
$Q_{LieJ}$ is not necessarily a homological vector field since $Q_{LieJ}^{2}$ is not necessarily zero without any condition for ${\underline{J}}{}$. In order to realize the condition (11) of a compatible $E$-$n$-form $J$, we modify $Q_{LieJ}$ by adding a term induced from a (pre)-$n$-plectic form $\omega=H$ asIn this section and the next section, a pre-$n$-plectic form on a (normal) base manifold $M$ is denoted by $H$ not $\omega$ to distinguish with ${\omega}_{grad}$.
${Q_{\mathrm{comp}}}$ =$={\{{{-\Theta_{Lie}+(-1)^{n}{\underline{J}}{}},{-}}\}}+(-1)^{n(n+1)/2}({\omega}_{grad}^{\flat})^{-1}\iota_{\rho(q)}^{n}H$
$=\rho^{i}_{a}(x)q^{a}\frac{\partial}{\partial x^{i}}-\frac{1}{2}C_{bc}^{a}(x)q^{b}q^{c}\frac{\partial}{\partial q^{a}}$
$+\left((-1)^{n}\rho^{i}_{a}z_{i}-C_{ab}^{c}(x)q^{b}y_{c}+(-1)^{n}\frac{1}{(n-1)!}J_{ab_{2}\ldots b_{n}}q^{b_{2}}\ldots q^{b_{n}}\right)\frac{\partial}{\partial y_{a}}$

### v4

@@ 4 Examples
$\bigl{(}\mathrm{d}^{\nabla}+{}^{E}\mathrm{d}^{\nabla}\bigr{)}\mu=-\sum_{k=0}^{n}\iota_{\rho}^{n+1-k}{\omega}.$ (4.7)
Expanding the equation by form degree of both sides, equation (4.7) is the following $n$ equations,
$\mathrm{d}^{\nabla}\mu_{n-1}=-\iota_{\rho}^{1}{\omega},$
$\mathrm{d}^{\nabla}\mu_{n-2}+{}^{E}\mathrm{d}^{\nabla}\mu_{n-1}=-\iota_{\rho}^{2}{\omega},$
@@
$\mathrm{d}^{\nabla}\mu_{k-1}+{}^{E}\mathrm{d}^{\nabla}\mu_{k}=-\iota_{\rho}^{n+1-k}{\omega},$
@@
${}^{E}\mathrm{d}^{\nabla}\mu_{0}={}^{E}\mathrm{d}\mu_{0}=-\iota_{\rho}^{n+1}{\omega}.$ (4.8)
The final equation (4.8) is the condition that $J=\mu_{0}$ is a compatible $E$-$n$-form (3.1).
Proposition 4.16.
@@
5 Higher Dirac structures on $\bm{E\oplus\wedge^{n-1}E^{*}}$
@@
Let $(E,\rho,[-,-])$ be a Lie algebroid over $M$. Then, an almost Lie algebroid structure is induced on $E\oplus\wedge^{n-1}E^{*}$. Three operations are constructed from the Lie algebroid operations. They are a bilinear symmetric form ${({{-},{-}})}\colon\Gamma\bigl{(}E\oplus\wedge^{n-1}E^{*}\bigr{)}\times\Gamma\bigl{(}E\oplus\wedge^{n-1}E^{*}\bigr{)}\rightarrow\Gamma\bigl{(}\wedge^{n-2}E^{*}\bigr{)}$, a bundle map ${\widehat{\rho}}\colon E\oplus\wedge^{n-1}E^{*}\rightarrow TM$ called the anchor map, and a bilinear bracket ${[{{-},{-}}]_{D}}\colon\Gamma\bigl{(}E\oplus\wedge^{n-1}E^{*}\bigr{)}\times\Gamma\bigl{(}E\oplus\wedge^{n-1}E^{*}\bigr{)}\rightarrow\Gamma\bigl{(}E\oplus\wedge^{n-1}E^{*}\bigr{)}$ called a (higher) Dorfman bracket. Let $u+\alpha,v+\beta\in\Gamma\bigl{(}E\oplus\wedge^{n-1}E^{*}\bigr{)}$, where $u,v\in\Gamma(E)$ and $\alpha,\beta\in\Gamma\bigl{(}\wedge^{n-1}E^{*}\bigr{)}$. Then three operations are defined by
${({{u+\alpha},{v+\beta}})}=\iota_{u}\beta+\iota_{v}\alpha,\qquad{\widehat{\rho}}(e)f=\rho(u)f,$ (5.1)
${[{{u+\alpha},{v+\beta}}]_{D}}=[u,v]+{}^{E}{\mathcal{L}}_{u}\beta-\iota_{v}{}^{E}\mathrm{d}\alpha+\iota_{u}\iota_{v}\bigl{(}\iota_{\rho}^{n+1}\omega\bigr{)},$ (5.2)
where the interior product $\iota_{v}$ is the contraction of an element $v\in\Gamma(E)$ with $\Gamma(E^{*})$, $\rho(u)$ is the anchor map of $E$, and ${}^{E}{\mathcal{L}}_{u}:=\iota_{u}{}^{E}\mathrm{d}+{}^{E}\mathrm{d}\iota_{u}$ is the $E$-Lie derivative. The Dorfman bracket (5.2) is not skew-symmetric but satisfies the Jacobi type identity,
${[{{{[{{u+\alpha},{v+\beta}}]_{D}}},{w+\gamma}}]_{D}}+{[{{{[{{v+\beta},{w+\gamma}}]_{D}}},{u+\alpha}}]_{D}}$
$\qquad+{[{{{[{{w+\gamma},{u+\alpha}}]_{D}}},{v+\beta}}]_{D}}=0.$ (5.3)
An algebroid satisfying equation (5.3) is called a Leibniz algebroid. Especially, we call
$\bigl{(}E\oplus\wedge^{n-1}E^{*},{({{-},{-}})},\rho,{[{{-},{-}}]_{D}}\bigr{)}$
a Lie $n$-algebroid, or a Vinogradov algebroid induced from a Lie algebroid $E$ [20]. We only consider this specific Lie $n$-algebroid in this paper. There are many analysis about general theories for Lie $n$-algebroids. In general, Lie $n$-algebroids are induced from a QP-manifold of degree $n$. A general concept of Lie $n$-algebroids is referred to, for instance, [7, 16, 17, 21, 29, 39, 41, 45].
Equations (5.1)–(5.2) does not contain $J$. The compatible condition (3.1), ${}^{E}\mathrm{d}J=-\iota_{\rho}^{n+1}\omega$, is described as an extra structure on the Lie $n$-algebroid, a higher Dirac structure[4, 8, 45].
Definition 5.1.
@@
$\Gamma(L)$ =$=\big{\{}u+{({{J},{u}})}\in\Gamma\bigl{(}E\oplus\wedge^{n-1}E^{*}\bigr{)}\mid u\in\Gamma(E)\big{\}}.$
Theorem 5.2.
@@
Moreover, using (5.2), the Dorfman bracket is computed as
@@
$=[u,v]+{}^{E}{\mathcal{L}}_{u}\iota_{v}J-\iota_{v}{}^{E}\mathrm{d}\iota_{u}J+\iota_{u}\iota_{v}\bigl{(}\iota_{\rho}^{n+1}\omega\bigr{)}$
$=[u,v]+{}^{E}{\mathcal{L}}_{u}\iota_{v}J-\iota_{v}{}^{E}{\mathcal{L}}_{u}J-\iota_{v}\iota_{u}{}^{E}\mathrm{d}J+\iota_{u}\iota_{v}\bigl{(}\iota_{\rho}^{n+1}\omega\bigr{)}.$ (5.4)
Using the formula ${}^{E}{\mathcal{L}}_{u}\iota_{v}-\iota_{v}{}^{E}{\mathcal{L}}_{u}=\iota_{[u,v]}$ analogous to the Cartan formula, (5.4) becomes
$=[u,v]+{({{J},{[u,v]}})}+\iota_{u}\iota_{v}{}^{E}\mathrm{d}J+\iota_{u}\iota_{v}\bigl{(}\iota_{\rho}^{n+1}\omega\bigr{)},$ (5.5)
which is an element of $\Gamma(L)$ again if the identity ${}^{E}\mathrm{d}J=-\iota_{\rho}^{n+1}\omega$ is satisfied. Note that we use $\iota_{u}\iota_{v}{}^{E}\mathrm{d}J=-\iota_{v}\iota_{u}{}^{E}\mathrm{d}J$. Since then (5.5) becomes
${[{{u+{({{J},{u}})}},{v+{({{J},{v}})}}}]_{D}}$ $=[u,v]+{({{J},{[u,v]}})}$
@@
${({{\alpha+u},{\beta+v}})}=\iota_{u}\beta+\iota_{v}\alpha,\qquad{\widehat{\rho}}(\alpha+u)f=\pi^{\sharp}(\alpha)(f),$
${[{{\alpha+u},{\beta+v}}]_{D}}=[\alpha,\beta]_{\pi,H}+{\mathcal{L}}^{\pi,H}_{\alpha}v-\iota_{\beta}\mathrm{d}_{\pi,H}u+\bigl{\langle}\otimes^{3}\pi,H\bigr{\rangle}(\alpha,\beta),$
@@
$[\alpha,\beta]_{\pi,H}={\mathcal{L}}_{\pi^{\sharp}(\alpha)}\beta-{\mathcal{L}}_{\pi^{\sharp}(\beta)}\alpha-\mathrm{d}(\pi(\alpha,\beta))+\iota_{\pi^{\sharp}(\alpha)}\iota_{\pi^{\sharp}(\beta)}H,\qquad\mathrm{d}_{\pi}u=[\pi,u]_{S},$
$\mathrm{d}_{\pi,H}u=\mathrm{d}_{\pi}u+\big{\langle}\bigl{(}\otimes^{2}\pi\otimes 1\bigr{)}H,u\big{\rangle},\qquad L^{\pi,H}_{\alpha}v=(\iota_{\alpha}\mathrm{d}_{\pi,H}+\mathrm{d}_{\pi,H}\iota_{\alpha})v.$
Let us introduce a bivector field $J\in\Gamma\bigl{(}\wedge^{2}TM\bigr{)}$ and define a subbundle defined by
$L$ =$=\big{\{}\alpha+{({{J},{\alpha}})}\in\Gamma(T^{*}M\oplus TM)\mid\alpha\in\Omega^{1}(M)\big{\}}.$
Iff $L$ is a Dirac structure, $J$ is a compatible $T^{*}M$-$2$-form, i.e., a bivector field satisfying ${}^{E}\mathrm{d}J=-\iota_{\pi}^{2}H$. This case is special. We can take $J=\pi$ since both are elements of $\Gamma\bigl{(}\wedge^{2}TM\bigr{)}$, and ${{}^{E}\mathrm{d}\pi=-\iota_{\pi}^{2}H}$ is equivalent to the condition that $(\pi,H)$ is a twisted Poisson structure.
@@ 6 Q-manifold description
Definition 6.1.
If an N-manifold ${\mathcal{M}}$ has a vector field $Q$ of degree $+1$ satisfying $Q^{2}=0$, it is called a Q-manifold, or a differential graded ((((dg)))) manifold.
@@
$Q$ $=\rho^{i}_{a}(x)q^{a}\frac{\partial}{\partial x^{i}}-\frac{1}{2}C_{bc}^{a}(x)q^{b}q^{c}\frac{\partial}{\partial q^{a}},$ (6.1)
where $\rho^{i}_{a}(x)$ and $C^{a}_{bc}(x)$ are local functions. We define a bundle map $\rho\colon E\rightarrow TM$ and the Lie bracket $[-,-]\colon\Gamma(E)\times\Gamma(E)\rightarrow\Gamma(E)$ as $\rho(e_{a}):=\rho^{i}_{a}(x)\partial_{i}$, $[e_{a},e_{b}]:=C_{ab}^{c}(x)e_{c}$, for the basis of sections $e_{a}\in\Gamma(E)$. Then, $Q^{2}=0$ is equivalent that the two operations $(\rho,[-,-])$ satisfy the definition of a Lie algebroid on $E$: all the identities of the anchor map and the Lie bracket in the Lie algebroid $E$ are obtained from the $Q^{2}=0$ condition.
@@
Definition 6.2.
If an N-manifold ${\mathcal{M}}$ has a graded symplectic form ${\omega}_{\rm grad}$ of degree $n$ and a vector field $Q$ of degree $+1$ satisfies $Q^{2}=0$, and ${\mathcal{L}}_{Q}{\omega}_{\rm grad}=0$, $({\mathcal{M}},{\omega}_{\rm grad},Q)$ is called a QP-manifold of degree $n$.
In a QP-manifold, if $n\neq 0$, there exists a Hamiltonian function $\Theta$ such that $Q={\{{{\Theta},{-}}\}}$. $Q^{2}=0$ is equivalent to ${\{{{\Theta},{\Theta}}\}}=0$.
For a Lie algebroid, we take an $n-1$-shifted cotangent bundle ${\mathcal{M}}=T^{*}[n-1]E[1]$ with a nonnegative integer $n-1$. Take local coordinates on $T^{*}[n-1]E[1]$, $(x^{i},q^{a},z_{i},y_{a})$ of degree $(0,1,n-1,n-2)$. Since $T^{*}[n-1]E[1]$ is a (graded) cotangent bundle, it has a canonical graded symplectic form of degree $n-1$, ${\omega}_{\rm grad}=\delta x^{i}\wedge\delta z_{i}+\delta q^{a}\wedge\delta y_{a}$, where $\delta$ is the graded de Rham differential on the graded manifold. A graded Poisson bracket for ${\omega}_{\rm grad}$ is defined by
${\{{{f},{g}}\}}=(-1)^{|f|+n}\iota_{X_{f}}\iota_{X_{g}}{\omega}_{\rm grad}.$
@@
$\big{\{}x^{i},z_{j}\big{\}}$ $=\delta^{i}_{j},\qquad\big{\{}z_{j},x^{i}\big{\}}=-\delta^{i}_{j},\qquad\{q^{a},y_{b}\}=\delta^{a}_{b},\qquad\{y_{b},q^{a}\}=-(-1)^{n-2}\delta^{a}_{b}.$
@@
$\Theta_{\rm Lie}$ $=\langle\rho(q),\,z\rangle+\frac{(-1)^{n-1}}{2}{({{[q,q]},{y}})}=(-1)^{n-1}\rho^{i}_{a}(x)z_{i}q^{a}+\frac{(-1)^{n-1}}{2}C_{bc}^{a}(x)q^{b}q^{c}y_{a}.$ (6.2)
Here $\langle$-,$\rangle$ - , - ⟩ is the pairing of $TM$ and $T^{*}M$ an (,){({{-},{-}})}( - , - ) is the pairing of $E$ and $E^{*}$. ${\{\Theta_{\rm Lie},\Theta_{\rm Lie}\}=0}$ is equivalent to $\rho(e_{a})=\rho^{i}_{a}(x)\partial_{i}$ and $C_{bc}^{a}(x)$ are the anchor map and the structure function of the Lie bracket of a Lie algebroid.
We can prove that $\iota_{Q_{\rm Lie}}{\omega}_{\rm grad}=-\delta\Theta_{\rm Lie}$, equivalently, $Q_{\rm Lie}=-\{\Theta_{\rm Lie},-\}$ in (6.2) satisfies ${Q_{\rm Lie}^{2}=0}$ and ${\mathcal{L}}_{Q_{\rm Lie}}{\omega}_{\rm grad}=0$ if and only if $E$ is a Lie algebroid. The local coordinate expression is
${Q_{\mathrm{Lie}}}={}$$-\{\Theta_{\rm Lie},-\}=\rho^{i}_{a}(x)q^{a}\frac{\partial}{\partial x^{i}}-\frac{1}{2}C_{bc}^{a}(x)q^{b}q^{c}\frac{\partial}{\partial q^{a}}+\bigl{(}(-1)^{n}\rho^{i}_{a}z_{i}-C_{ab}^{c}(x)q^{b}y_{c}\bigr{)}\frac{\partial}{\partial y_{a}}$
$-\left(\partial_{i}\rho^{j}_{a}z_{j}q^{a}-\frac{(-1)^{n}}{2}\partial_{i}C_{bc}^{a}(x)q^{b}q^{c}y_{a}\right)\frac{\partial}{\partial z_{i}}.$
Here $\partial_{i}=\frac{\partial}{\partial x^{i}}$. ${Q_{\mathrm{Lie}}}$ is also denoted by ${\mathcal{L}}_{Q}={Q_{\mathrm{Lie}}}$ and called the tangent lift of the vector field $Q$ in (6.1). ${Q_{\mathrm{Lie}}}$ is the homological vector field such that $Q={Q_{\mathrm{Lie}}}|_{E[1]}$. $T^{*}[n-1]E[1]$ has a QP-manifold structure induced from a Lie algebroid structure on $E$.
@@
${\underline{J}}{}$ $:=\frac{1}{n!}J_{a_{1}\dots a_{n}}(x)q^{a_{1}}\cdots q^{a_{n}},$
@@
$\frac{1}{n!}J_{a_{1}\dots a_{n}}(x)$ $:=J(e_{a_{1}},\dots,e_{a_{n}}).$
The term ${\underline{J}}{}$ is summed to $\Theta_{\rm Lie}$ as $\Theta_{{\rm Lie}\,J}=\Theta_{\rm Lie}+(-1)^{n-1}{\underline{J}}{}$, and consider the vector field of degree $+1$
$Q_{{\rm Lie}\,J}$ =$=-{\{{{\Theta_{{\rm Lie}\,J}},{-}}\}}.$
$Q_{{\rm Lie}\,J}$ is not necessarily a homological vector field since $Q_{{\rm Lie}\,J}^{2}$ is not necessarily zero without any condition for ${\underline{J}}{}$. In order to realize the condition (3.1) of a compatible $E$-$n$-form $J$, we modify $Q_{{\rm Lie}\,J}$ by adding a term induced from a (pre)-$n$-plectic form $\omega=H$ as [^5] In this section and the next section, a pre-$n$-plectic form on a (normal) base manifold $M$ is denoted by $H$ not $\omega$ to distinguish with ${\omega}_{\rm grad}$.
${Q_{\mathrm{comp}}}={}$${\{{{-\Theta_{\rm Lie}+(-1)^{n}{\underline{J}}{}},{-}}\}}+(-1)^{n(n+1)/2}\bigl{(}{\omega}_{\rm grad}^{\flat}\bigr{)}^{-1}\iota_{\rho(q)}^{n}H$
[…]
$+\left((-1)^{n}\rho^{i}_{a}z_{i}-C_{ab}^{c}(x)q^{b}y_{c}+(-1)^{n}\frac{1}{(n-1)!}J_{ab_{2}\dots b_{n}}q^{b_{2}}\cdots q^{b_{n}}\right)\frac{\partial}{\partial y_{a}}$